OpenStudy (mathmath333):

the ratio of factorial of a number x to the square of the factorial of another number,which when increased by 50% gives the required number,is 1.25. find the number x?

OpenStudy (ikram002p):

p! / (q!)^2 so u mean q+0.5 q =p ?

OpenStudy (mathmath333):

optinons a. 6 b 5 c 9

OpenStudy (mathmath333):

well the interpretation of que is itself confusing to me

OpenStudy (ikram002p):

ohkk , so its like this :) x! / (q!)^2 + 50 % x! / (q!)^2 = 1.25

OpenStudy (ikram002p):

@ganeshie8 is it correct so far ?

OpenStudy (mathmath333):

the ans shows 6!/(4!^2)=1.25, but i cant get it

OpenStudy (ikram002p):

im still trying to interpt it

OpenStudy (ikram002p):

btw in the question which referse to what xD

OpenStudy (mathmath333):

x refers to 6

OpenStudy (ikram002p):

this which http://prntscr.com/4af8q0

OpenStudy (mathmath333):

i think que may be wrong

OpenStudy (anonymous):

4+50% of 4 = 6

OpenStudy (ikram002p):

ok ok its like this lol x!/q! =1.25 x=q+0.5 q now it make sense :)

OpenStudy (anonymous):

so its, q+50% of q = x

OpenStudy (ikram002p):

so q!/(q+0.5q )! ^2 =1.25 solve for q

OpenStudy (anonymous):

(q+0.5q)!/(q!)^2 =1.25

OpenStudy (ikram002p):

oh sorry typo sau

OpenStudy (ikram002p):

so q=x/1.5 (q+0.5q)!/(q!)^2 =1.25 (1.5 ) q! / q!^2 =1.25 1.5 =1.25 q!

OpenStudy (ikram002p):

mmm

OpenStudy (anonymous):

that last step cannot be done

ganeshie8 (ganeshie8):

i think your interpretation is correct, 150% of q must give the x x = 1.5q

OpenStudy (anonymous):

its , (1.5q)!/(q!)^2 =1.25 q! cannot be canceled

OpenStudy (ikram002p):

yeah mmm being ratio 5/6 still dont know

OpenStudy (ikram002p):

lol ok xD i thought this thing is correct xD (ab )! = a (b!) = a! (b)

OpenStudy (ikram002p):

ok we can interept it like this 1.5 q should be integer

OpenStudy (ikram002p):

also q should be integer

ganeshie8 (ganeshie8):

and only 6 makes it an integer is it

OpenStudy (ikram002p):

we should reason that

ganeshie8 (ganeshie8):

ofcourse out of the given options

OpenStudy (ikram002p):

lets see if there is another way

ganeshie8 (ganeshie8):

that looks good to me, 5 and 9 are odd so they cannot make 1.5q an integer... so we can conlude the answer is 6

OpenStudy (ikram002p):

(3/2 q ) then q is even hehe xD yeah options

OpenStudy (mathmath333):

thns

OpenStudy (ikram002p):

@ganeshie8 how to go far from here ?  ok lol what i wanna say is n=1,2,3,4,..... so check whick n satisfy :3 that would be n=2 q=2n=4 thus x=6